# Comparing Left and Right Quotient Sets in Groups

### v2

@@ 2.3. Cardinality of AA
Let $G$ be a group with $A\subset G$ and $|A|=n$. Then $D_{A}$ contains no connected component (other then the diagonal) with more than $n$ elements.
@@
Suppose, for contradiction, $C$ was a connected component with more than $n$ elements. By the pigeonhole principle, two vertices in $C$ have the same first coordinate. Since $C$ is a clique by 2.1 (4), this contradicts 2.1 (2b). ∎
@@
Let $G$ be a group with $A\subset G$ and $|A|=4$. If $G$ does not have an element of order 2 and the largest possible clique in $D_{A}$ is $K_{4}$, the number of connected components in $D_{A}$ is equal to the number of connected components in $D_{A^{-1}}$.
@@
Consider a triangle $\triangle=(\alpha_{1},\beta_{1})(\alpha_{2},\beta_{2})(\alpha_{3},\beta_{3})$. By 2.1 (2b), we know that the same number can appear at most twice in the coordinates $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$, $\alpha_{3}$, and $\beta_{3}$. Moreover the same number appears at most once for an $\alpha_{i}$ and at most once for a $\beta_{j}$, where $1\leq i,j\leq 3$. Therefore, there are two cases to consider.
Case 1: There are 4 distinct elements in $\triangle$. Because of the condition that there is no element of order 2 in $G$, there cannot be an edge between $(i,j)$ and $(j,i).$ Therefore, up to relabeling, $\triangle=(1,2)(2,3)(3,4)$. Then, $D_{A}$ and $D_{A^{-1}}$ look as follows.
@@
Therefore, in this case, the number of connected components stays the same. Additionally there may be a triangle $\triangle^{\prime}=(3,1)(4,2)(1,4)$ and its reflection under $T$. However, we can check that the image of this graph under $\phi$ is itself as shown below.
@@
Case 2: There are 3 distinct elements in $\triangle$. The only possibility in this case is that up to relabelling, $\triangle=(1,2)(2,3)(3,1).$ Notice that this triangle cannot be part of a $K_{4}$ because $(4,4)$ is not in the same connected component as $\triangle$ and every other coordinates will cause contradiction to Lemma 2.1 (2b). Then, $D_{A}$ and $D_{A^{-1}}$ are as follows.
@@
In the second case, we see it is impossible to place a second triangle without two vertices being on the same axis, violating 2.1 (2b). ∎
@@
Assume $D_{A}$ has a non-diagonal clique of size 3 or greater. This is impossible if $|A|\in\left\{1,2\right\}$ by 2.7. If $|A|=3$, then the only possible triangles are $(1,2)(3,1)(2,3)$ and $(2,1)(1,3)(3,2)$. Since any $D_{A}$ is invariant under transposition by $T$, if $D_{A}$ has a triangle, then it has both. Any other edge would either (1) connect the triangles to the diagonal, which is forbidden, or (2) connect the triangle to each other, which creates a component with more $3$ vertices, contradicting 2.7. Thus, with $|A|=3$, $D_{A}$ and $D_{A^{-1}}$ must have the same component counts. So we have shown for any group $G$ if $|A|\leq 3$ then $|AA^{-1}|=|A^{-1}A|$.
Notice how the above proof does not use 2.1 (2c), so it holds for any group, yielding the first part of the theorem.
Next, suppose that $G$ has no elements of order $2$. Then note that the size of any clique in $D_{A}$ is at most 4 by 2.7. We claim $D_{A}$ cannot have a $K_{4}$ other than the diagonal. Suppose $D_{A}$ did have such a $K_{4}$. Let the vertices of this $K_{4}$ be $(1,j_{1})$, $(2,j_{2})$, $(3,j_{3})$, $(4,j_{4})$. Up to relabeling, we may assume $j_{1}=2$. Further, we may assume up to relabeling that $j_{2}=3$. This forces $j_{4}\in\left\{1,4\right\}$, which implies it equals $1$. Therefore, $j_{3}=4$. This clique is impossible if $G$ has no elements of order $2$. Indeed,
$a_{1}a_{3}^{-1}~=~(a_{1}a_{2}^{-1})(a_{2}a_{3}^{-1})~=~(a_{3}a_{4}^{-1})(a_{4}a_{1}^{-1})~=~a_{3}a_{1}^{-1}.$
Therefore, $D_{A}$’s components are $K_{3}$’s and $K_{2}$’s. 2.8 implies that $\phi$ induces a bijection on the $K_{3}$’s, and hence on the $K_{2}$’s as well, by an edge counting argument combined with the fact that there are no $K_{4}$’s. Hence, we have shown if $G$ is a group with no elements of order 2 and $|A|\leq 4$, then $|AA^{-1}|=|A^{-1}A|$. ∎
@@
If $G$ has elements of order $2$, then we may no longer use 2.1 (2c) in the previous statements. Indeed, we check with Sage [The25] that the set
@@
3. Future Work
In a finite subset $S$ of a group $G$, one may ask how many finite subsets $A\subseteq S$ have the left quotient set larger than the right quotient set. If $S$ is symmetric (i.e., $w\in S$ if and only if $w^{-1}\in S$), this is equal to the number of finite subsets with right quotient set larger than left quotient set. Indeed, if $|A^{-1}A|>|AA^{-1}|$, then we can replace $A$ with $B\coloneqq A^{-1}$ to get $|B^{-1}B|<|BB^{-1}|$, and vice-versa. In particular, this implies that the quantity
@@ 3. Future Work
where $A\subseteq S$ is chosen by including elements at random with probability $1/2$.
On the other hand, the second moment (i.e., the variance)
@@
is nonzero for groups where $|AA^{-1}|\neq|A^{-1}A|$ for some finite subset $A\subseteq S$. If $A$ is a subgroup then trivially one has the two cardinalities are the same, as in that case $A=A^{-1}$. Having to avoid such special cases is common; for example, there is non-Benford behavior if the support of certain random variables live in a coset of a group (see [MN08], in particular the discussion related to a result of Levy [Lév39]).
We can cast this problem specifically to the free group.
@@
Let $B_{N}\subseteq F_{2}$ be the set of words of length $\leq N$. With respect to the uniform probability measure on the subsets of $B_{N}$, what is $\mathrm{Var}(|A^{-1}A|-|AA^{-1}|)$?
A technique for studying the moments of $|AA^{-1}|$ and related quantities is the graph theoretic framework in [LMO13]. This paper suggests that to study $\mathbb{P}(w_{1},\dots,w_{k}\notin AA^{-1})$, we should look at the graph with vertex set $S$ and edge relation $u\sim v$ if $uv^{-1}=w_{m}$ or $vu^{-1}=w_{m}$ for some $m=1,\dots,k$. It would be interesting to find a structure which encapsulates both the the above graphs and the graphs $D_{A}$ and $D_{A^{-1}}$ which we introduced.
There are still other unanswered questions about the values $|AA^{-1}|-|A^{-1}A|$ can take in various groups.
Question 3.2 (Answered in [MS25] and [HKLM14]).
@@
Of course, $G$ being non-abelian is a necessary condition for $|A^{-1}A|\neq|AA^{-1}|$. However, one can check that $|AA^{-1}|=|A^{-1}A|$ for all subsets $A$ of the symmetric group $S_{3}$. Hence, being non-abelian is not a sufficient condition. It is possible that $|AA^{-1}|=|A^{-1}A|$ is true for all $A\subseteq S_{3}$ only because $S_{3}$ is sufficiently small [^3] As $|S_{n}|=n!$ , the number of subsets is $2^{n!}$, which grows very rapidly; $2^{3!}=64$, $2^{4!}$ is approximately $1.7\cdot 10^{7}$, while $2^{5!}$ exceeds $10^{36}$; straightforward sampling cannot hope but to explore a very small percentage of possible $A$’s. [^4] A relevant fact here is that $|AA|<2|A|$ is a sufficient but not necessary condition for $|AA^{-1}|=|A^{-1}A|$ (see [Tao]). One consequence of this is that no set $A$ of size $|A|>|G|/2$ can have $|AA^{-1}|\neq|A^{-1}A|$. Since $|S_{3}|=6$, this means all subsets of $S_{3}$ where $|AA^{-1}|\neq|A^{-1}A|$ have size $|A|>|S_{3}|/2$ or $|A|\leq 3$, and by 1.3, this is enough to prove $|AA^{-1}|=|A^{-1}A|$ for all subsets $A\subseteq S_{3}$ only in terms of the sizes of subsets.. Therefore, to find a more robust example, we ask whether there are infinitely many non-abelian groups where $|A^{-1}A|\neq|AA^{-1}|$ is impossible.
Question 3.3 (Answered in [MS25] and [HKLM14]).
@@
Remark 3.4.
Following the initial draft of this paper, the sixth author answered questions 3.2. and 3.3. in the preprint [MS25]. Shortly following this preprint, the authors of the present paper were directed to [HKLM14] which answered the questions previously and independently. Both answer the question by providing a classification of groups $G$ for which $|AA^{-1}|=|A^{-1}A|$ for all $A\subseteq G$. Question 3.3. is answered affirmatively with the Hamiltonian $2$-groups $Q_{8}\times(C_{2})^{n}$ being the unique infinite family of groups where $|AA^{-1}|=|A^{-1}A|$ holds for all subsets. This classification appears as [HKLM14, Theorem 7.4.] for all finite and infinite groups and as [MS25, Theorem 1.1.] where the classification is only for finite groups. We are grateful to Liubomir Chiriac for pointing us to [HKLM14] after reading a previous draft of this paper.
An alternative way to characterize groups where $|AA^{-1}|\neq|A^{-1}A|$ is possible would be to find a finite collection of difference graphs which serve as minimal counterexamples to the statement $|AA^{-1}|=|A^{-1}A|$ for all $A\subseteq G$. We conjecture that such a collection exists. That is to say, there should be a finite collection of graphs $\{H_{1},\dots,H_{k}\}$ such that, if $|AA^{-1}|\neq|A^{-1}A|$ for some $A\subseteq G$, then there is some $B\subseteq G$ such that $D_{B}\cong H_{i}$ for some $i=1,\dots,k$. A counterexample to our conjecture would be an infinite sequence of groups $G_{\ell}$ such that for any $N>0$, there is some group $G_{m}$ in the collection for which $|AA^{-1}|=|A^{-1}A|$ for all sets $A$ of size <N<N, but $|AA^{-1}|\neq|A^{-1}A|$ for some $A$ of size $\geq N$.

### v3

@@ 2.3 Cardinality of AA
Let $G$ be a group with $A\subset G$ and $|A|=n$. Then $D_{A}$ contains no connected component with more than $n$ elements.
@@
Suppose, for contradiction, $C$ was a connected component with more than $n$ elements. By the pigeonhole principle, two vertices in $C$ have the same first coordinate. Since $C$ is a clique by Lemma 2.1 (4), this contradicts Lemma 2.1 (2b). ∎
@@
Let $G$ be a group with $A\subseteq G$ and $|A|=4$. If $G$ does not have an element of order 2, then the number of connected components in $D_{A}$ is equal to the number of connected components in $D_{A^{-1}}$.
@@
Suppose $D_{A}$ contains a $K_{4}$ clique. Then, up to relabeling and transposition, this clique has the form $(1,2)(2,3)(3,4)(4,1)$. Indeed, given a $K_{4}$ of the form $(1,j_{1})(2,j_{2})(3,j_{3})(4,j_{4})$, we may assume $j_{1}=2$ and $j_{2}=3$ up to relabeling. This forces $j_{4}=1$ and $j_{3}=4$. This clique is impossible if $G$ has no elements of order $2$ since $a_{1}a_{3}^{-1}=(a_{1}a_{2}^{-1})(a_{2}a_{3}^{-1})=(a_{3}a_{4}^{-1})(a_{4}a_{1}^{-1})=a_{3}a_{1}^{-1}$ and this contradicts Lemma 2.1 (2c) given the assumption that $G$ has no elements of order $2$.
It remains to consider the case where the largest clique is a $K_{3}$, which we call a triangle. Consider a triangle $\triangle=(\alpha_{1},\beta_{1})(\alpha_{2},\beta_{2})(\alpha_{3},\beta_{3})$. By Lemma 2.1 (2b), we know that the same number can appear at most twice in the coordinates $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$, $\alpha_{3}$, and $\beta_{3}$. Moreover the same number appears at most once for an $\alpha_{i}$ and at most once for a $\beta_{j}$, where $1\leq i,j\leq 3$. Therefore, there are two cases to consider.
Case 1: There are 4 distinct elements in $\triangle$. Because of the condition that there is no element of order 2 in $G$, there cannot be an edge between $(i,j)$ and $(j,i).$ Therefore, up to relabeling, $\triangle=(1,2)(2,3)(3,4)$. By elementary group operations, the edge $[(1,3),(2,4)]$ and its transpose are also in $D_{A}$. Hence, $D_{A}$ and $D_{A^{-1}}$ look as in Figure 2. Therefore, in this case, the number of connected components stays the same.
@@
Figure 2: Up to relabeling, $\triangle=(1,2)(2,3)(3,4)$ induces $D_{A}$ and $D_{A^{-1}}$ above.
Additionally there may be a triangle $\triangle^{\prime}=(3,1)(4,2)(1,4)$ and its reflection under $T$. However, we can check that the image of this graph under $\phi$ is itself as in Figure 3 below.
@@
Figure 3: Up to relabeling, $\triangle^{\prime}=(3,1)(4,2)(1,4)$ induces $D_{A}$ and $D_{A^{-1}}$ above.
Case 2: There are 3 distinct elements in $\triangle$. The only possibility in this case is that up to relabeling, $\triangle=(1,2)(2,3)(3,1).$ Notice that this triangle cannot be part of a $K_{4}$ because $(4,4)$ is not in the same connected component as $\triangle$ and all other coordinates will cause a contradiction with Lemma 2.1 (2b). Then, $D_{A}$ and $D_{A^{-1}}$ are as in Figure 4 below.
@@
Figure 4: Up to relabeling, $\triangle=(1,2)(2,3)(3,1)$ induces $D_{A}$ and $D_{A^{-1}}$ above.
In the second case, we see it is impossible to place a second triangle without two vertices being on the same axis, violating Lemma 2.1 (2b). ∎
@@
Assume $D_{A}$ has a non-diagonal clique of size 3 or greater. This is impossible if $|A|\in\left\{1,2\right\}$ by Lemma 2.8. If $|A|=3$, then the only possible triangles are $(1,2)(3,1)(2,3)$ and $(2,1)(1,3)(3,2)$. Since any $D_{A}$ is invariant under transposition by $T$, if $D_{A}$ has a triangle, then it has both. Any other edge would either (1) connect the triangles to the diagonal, which is forbidden, or (2) connect the triangles to each other, which creates a component with more $3$ vertices, contradicting Lemma 2.8. Thus, with $|A|=3$, $D_{A}$ and $D_{A^{-1}}$ must have the same component counts. So, we have shown for any group $G$ if $|A|\leq 3$ then $|AA^{-1}|=|A^{-1}A|$.
Notice how the above proof does not use Lemma 2.1 (2c), so it holds for any group, yielding the first part of the theorem.
Next, suppose that $G$ has no elements of order $2$. If $|A|=4$, then Lemma 2.9 shows that $\phi$ preserves the number of connected components. Hence, we have shown if $G$ is a group with no elements of order 2 and $|A|\leq 4$, then $|AA^{-1}|=|A^{-1}A|$. ∎
@@
If $G$ has elements of order $2$, then we may no longer use Lemma 2.1 (2c) and proof of Lemma 2.9 fails. Indeed, we may check with Sage [14] that the set
@@
§3  Future Work
In a finite subset $S$ of a group $G$, one may ask how many finite subsets $A\subseteq S$ have the left quotient set larger than the right quotient set. If $S$ is symmetric (i.e., $w\in S$ if and only if $w^{-1}\in S$), the number of subsets with the left quotient set larger than the right equals the number of subsets with the right quotient set larger than the left. Indeed, if $|A^{-1}A|>|AA^{-1}|$, then we can replace $A$ with $B\coloneqq A^{-1}$ to get $|B^{-1}B|<|BB^{-1}|$, and vice-versa. In particular, this implies that the quantity
@@
where $A\subseteq S$ is chosen by including elements at random with probability $1/2$. On the other hand, the variance
@@
is nonzero for groups where $|AA^{-1}|\neq|A^{-1}A|$ for some finite subset $A\subseteq S$. This suggests the following question in the free group $F_{2}$.
@@
Let $B_{N}\subseteq F_{2}$ be the set of words of length no greater than $N$. With respect to the uniform probability measure on the subsets of $B_{N}$, what is $\mathrm{Var}(|A^{-1}A|-|AA^{-1}|)$?
When studying the moments of $|AA^{-1}|$, one must handle dependence between the event $x\in AA^{-1}$ and the event $y\in AA^{-1}$. A technique for handling this dependence is via constructing the graph $C$ with vertex set $V(C)=S$ and edge set $E(C)=\{(u,v)\in S\times S\mid uv^{-1}=w_{m}\text{ or }vu^{-1}=w_{m}\text{ for some }m=1,\dots,k\}$. Then, $\mathbb{P}(w_{1},\dots,w_{k}\notin AA^{-1})$ is the probability that $A$ is chosen to be a vertex cover of $C$. This technique is described in [7] for finite sets of integers.
There are other unanswered questions about the values $|AA^{-1}|-|A^{-1}A|$ can take in various groups.
Question 3.2 (Answered in [11] and [5]).
@@
Of course, $G$ being non-abelian is a necessary condition for $|A^{-1}A|\neq|AA^{-1}|$. However, one can check that $|AA^{-1}|=|A^{-1}A|$ for all subsets $A$ of the symmetric group $S_{3}$. Hence, being non-abelian is not a sufficient condition.
Question 3.3 (Answered in [11] and [5]).
@@
Following the initial draft of this paper, the sixth author answered Questions 3.2 and 3.3 in the preprint [11]. Shortly following this preprint, the authors of the present paper were directed to [5] which answered the questions previously and independently. Both answer the question by providing a classification of groups $G$ for which $|AA^{-1}|=|A^{-1}A|$ for all $A\subseteq G$. Question 3.3 is answered affirmatively with the Hamiltonian $2$-groups $Q_{8}\times(C_{2})^{n}$ being the unique infinite family of groups where $|AA^{-1}|=|A^{-1}A|$ holds for all subsets.
This classification appears as [5, Theorem 7.4.] for all finite and infinite groups and as [11, Theorem 1.1.] where the classification is only for finite groups. We are grateful to Liubomir Chiriac for pointing us to [5] after reading a previous draft of this paper. Given that Questions 3.2 and 3.3 are solved, we are interested in the following extension. Define the family of functions on groups $\{f_{n}\}_{n\geq 1}$ given by $f_{n}(G)=\sup_{A\subseteq G,|A|\leq n}(|AA^{-1}|-|A^{-1}A|)$.
Question 3.4.
Fix $c\geq 0$. What are the necessary and sufficient conditions on a group $G$ such that $f_{n}(G)\leq c$ for $n\geq 0$?
Question 3.5.
For what groups $G$ does $f_{n}(G)\to\infty$ as $n\to\infty$?
The $c=0$ case of Question 3.4 is exactly Question 3.2 while the case $c>0$ asks for the maximal difference $|AA^{-1}|-|A^{-1}A|$ when $A\subseteq G$ is a subset of size $n$. We believe these questions could be a productive area for future research.
Acknowledgements.
The authors gratefully acknowledge the referee’s careful reading of the manuscript and their constructive comments. The authors also thank Carl Pomerance and Mel Nathanson for stimulating discussions; the genesis of this work came from conversations the fifth-named author had with them at the Integers Conference in Georgia in 2025. This project took place at the SMALL REU program at Williams College and was funded by the Finnerty Fund from Williams College and the National Science Foundation (Grant DMS2241623). The authors are grateful for the support of Amherst College, Princeton University, the University of Michigan, the University of Wisconsin, and Williams College.